# CUPID: Curating Data your Robot Loves with Influence Functions

### v1

@@ Abstract
In robot imitation learning, policy performance is tightly coupled with the quality and composition of the demonstration data. Yet, developing a precise understanding of how individual demonstrations contribute to downstream outcomes—such as closed-loop task success or failure—remains a persistent challenge. We propose Cupid, a robot data curation method based on a novel influence function-theoretic formulation for imitation learning policies. Given a set of evaluation rollouts, Cupid estimates the influence of each training demonstration on the policy’s expected return. This enables ranking and selection of demonstrations according to their impact on the policy’s closed-loop performance. We use Cupid to curate data by 1) filtering out training demonstrations that harm policy performance and 2) subselecting newly collected trajectories that will most improve the policy. Extensive simulated and hardware experiments show that our approach consistently identifies which data drives test-time performance. For example, training with less than 33% of curated data can yield state-of-the-art diffusion policies on the simulated RoboMimic benchmark, with similar gains observed in hardware. Furthermore, hardware experiments show that our method can identify robust strategies under distribution shift, isolate spurious correlations, and even enhance the post-training of generalist robot policies. Additional materials are made available at: https://cupid-curation.github.io. 00footnotetext: Correspondence to: cagia@cs.stanford.edu
@@ 2 Related Work
Data Curation in Robotics. Assembling larger and more diverse datasets has been central to scaling efforts in robot imitation learning [7, 8, 17, 18, 19, 20, 21], yet how to extract greater utility from these datasets remains an open question. Several works have explored data augmentation [22, 23, 24, 25, 26] and mixture optimization [27]. Only recently has attention shifted to valuating individual demonstrations for data curation [9, 10, 11]. Hejna et al. [10] estimate demonstration quality offline via mutual information—without considering policy performance. Closest to our work is Demo-SCORE [11], which trains classifiers to distinguish successful and failed rollouts across multiple policy checkpoints. In contrast, we directly measure the causal influence of each demonstration on the policy’s expected return, providing a signal that (a) does not require observing both successes and failures, (b) uses only a single policy checkpoint, (c) is robust to spurious correlations in the policy’s rollout distribution, and (d) naturally extends to selecting new data, whereas [10, 11] exclusively focus on filtering existing training sets.
@@ 3 Background: Data Attribution via Influence Functions
where $H_{\theta}=\frac{1}{n}\sum_{i=1}^{n}\nabla_{\theta}^{2}\ell(z^{i};\theta(\mathcal{D}))$ denotes the Hessian of the training loss111 To reduce the computational cost of Eq. 1, we use TRAK [2], which leverages random projections and a Gauss-Newton Hessian approximation for efficient influence estimation. This also makes the influence function amenable to the non-smooth, non-convex loss functions in practical deep learning problems, so we assume Eq. 1 is well-defined throughout this paper. [13]. Therefore, we can use the influence function to directly approximate the leave-one-out influence $\Psi^{\mathrm{out}}$ of a sample $z^{i}\in\mathcal{D}$ as $\Psi^{\mathrm{out}}_{\mathrm{inf}}(\hat{z},z^{i}):=-\frac{1}{n}\Psi_{\mathrm{inf}}(\hat{z},z^{i})$. In addition, for $z\not\in\mathcal{D}$ we similarly define the add-one-in influence as $\Psi^{\mathrm{in}}_{\mathrm{inf}}(\hat{z},z):=\frac{1}{n}\Psi_{\mathrm{inf}}(\hat{z},z)\approx f(\hat{z};\theta(\mathcal{D}\cup\{z\}))-f(\hat{z};\theta(\mathcal{D}))$ with $z$ excluded from the Hessian $H_{\theta}$.
@@ 4 Problem Formulation
Let $\Xi^{+}_{k}=\left\{S\subseteq\mathcal{H}\;\middle|\;|S|=k\right\}$ denote all possible $k$-demonstration subsets of a holdout dataset =$\mathcal{H}=\{\xi^{1},\ldots,\xi^{n^{\prime}}\}$, where $k\leq n^{\prime}$. Determine which $k$ demonstrations should be added to $\mathcal{D}$ from $\mathcal{H}$ to maximize policy performance with respect to the task objective $J$. That is, find
@@
[…]
@@ 5.1 Demonstration-Performance Influence
Assume that $\theta(\mathcal{D})=\arg\min_{\theta^{\prime}}\mathcal{L}_{\mathrm{bc}}(\theta^{\prime};\mathcal{D})$, that $\mathcal{L}_{\mathrm{bc}}$ is twice differentiable in $\theta$, and that $H_{\mathrm{bc}}\succ 0$ is positive definite (i.e., $\theta(\mathcal{D})$ is not a saddle point)\@footnotemark. Then, it holds that [^2] Note that the fraction $1/H$ appears from the assumption that all trajectories have equal length, which we make purely for notational simplicity without loss of generality. We refer to Section D.2 for the variable length case.
@@
[…]
@@
1 Input: Policy $\pi_{\theta}$, training data $\mathcal{D}$, demonstration $\xi$, data attribution method $\Psi$
[…]
@@ 9 Limitations
Selection methods. While the greedy selection procedures used in Eq. 4 and Eq. 5 are tractable to optimize and often improve over quality- and similarity-based measures [3], they ignore the interactions between demonstrations in the curated set [14, 47]. This can temper performance gains when the size of the curated set is large. Future work should investigate higher-order approximations that consider the joint diversity of the curated dataset, as is common in the active learning literature (e.g., [51, Sec. 4.3]).
@@ Acknowledgments
The appendix offers additional details w.r.t. the implementation of Cupid (Appendix A), the experiments conducted (Appendix B), along with extended results and analysis (Appendix C), and finally, supporting derivations for our data curation methods (Appendix D). Qualitative results are made available at: https://cupid-curation.github.io.
[…]
@@ A.1 Influence Functions for Diffusion Policies
Assume that $\theta(\mathcal{D})=\arg\min_{\theta^{\prime}}\mathcal{L}_{\mathrm{bc}}(\theta^{\prime};\mathcal{D})$, that $\mathcal{L}_{\mathrm{bc}}$ is twice differentiable in $\theta$, and that $H_{\mathrm{bc}}\succ 0$ is positive definite (i.e., $\theta(\mathcal{D})$ is not a saddle point)\@footnotemark. Then, it holds that
@@
$\ell(s,a;\pi_{\theta^{\prime}}):=\mathbb{E}_{\epsilon^{i},i}\left[||\epsilon^{i}-\epsilon_{\theta^{\prime}}(\sqrt{\bar{\alpha}_{i}}a+\sqrt{1-\bar{\alpha}_{i}}\epsilon^{i},s,i)||^{2}\right],$ (7)
@@
Influence Approximations: Since the denoising loss $\ell$ in Eq. 7 is proportionate to the variational lower bound on the action log-likelihood $\log\pi_{\theta}(a|s)$, it may seem intuitive to substitute $\nabla_{\theta}\log\pi_{\theta}(a^{\prime}|s^{\prime})$ with $-\nabla_{\theta}\ell(s^{\prime},a^{\prime};\pi_{\theta})$—assuming gradient alignment—to approximate the action influence (Eq. 2) as
@@
A similar approach is taken by Georgiev et al. [37] for attributing the generations of image-based diffusion models. However, consistent with more recent results in the data attribution literature [36, 54], we find this approximation to work poorly in practice, with highly influential training samples $(s,a)\in\mathcal{D}$ rarely reflecting the test-time transitions $(s^{\prime},a^{\prime})\in\tau$ over which the action influences are computed. Instead, we follow the approach of Zheng et al. [36], which entails replacing both $\log\pi_{\theta}(a^{\prime}|s^{\prime})$ and $\ell(s,a;\pi_{\theta})$ in Eq. 2 with a surrogate, label-agnostic output function $\ell_{\mathrm{square}}(s,a;\pi_{\theta}):=\mathbb{E}_{\epsilon^{i},i}[||\epsilon_{\theta}(\sqrt{\bar{\alpha}_{i}}a+\sqrt{1-\bar{\alpha}_{i}}\epsilon^{i},s,i)||^{2}]$, making our final approximation of the action influence
@@
    Using $\ell_{\mathrm{square}}$ significantly reduces the computational cost of computing action influences for policies with high-dimensional action spaces, because the $\ell^{2}$-norm collapses the model’s prediction into a scalar $||\epsilon_{\theta}(\sqrt{\bar{\alpha}_{i}}a+\sqrt{1-\bar{\alpha}_{i}}\epsilon^{i},s,i)||^{2}$. As a result, computing Eq. 9 requires only a single model gradient $\nabla_{\theta}\ell_{\mathrm{square}}$ per training and test sample. In contrast, while the technique proposed by Lin et al. [54] offers a more accurate estimate of the leave-one-out divergence $D_{\mathrm{KL}}(\pi_{\theta(\mathcal{D})}(a^{\prime}|s^{\prime})||\pi_{\theta(\mathcal{D}\setminus(s,a))}(a^{\prime}|s^{\prime}))$, its computational cost scales linearly with the dimensionality of the model’s output, which may be prohibitive.
@@ A.2 CUPID Hyperparameters
Action Influence (Eq. 9): The action influence requires computing the gradient of an expectation $\nabla_{\theta}\ell_{\mathrm{square}}(s,a;\pi_{\theta})=\nabla_{\theta}\mathbb{E}_{\epsilon^{i},i}[||\epsilon_{\theta}(\sqrt{\bar{\alpha}_{i}}a+\sqrt{1-\bar{\alpha}_{i}}\epsilon^{i},s,i)||^{2}]$. For all tasks, we approximate the expectation using a batch of $B=64$ samples $(\epsilon^{(b)},i^{(b)})$, where $\epsilon^{(b)}\sim\mathcal{N}(0,1)$ and $i^{(b)}\sim\mathcal{U}[0,T)$ are sampled independently.
@@ B.4 Baseline Details
Comparison to Performance Influence (Cupid): One can interpret Success Similarity as replacing the action influence $\Psi_{a\text{-}\mathrm{inf}}((s^{\prime},a^{\prime}),(s,a))$ (Eq. 2) with a state-based proxy $-D(\phi(s^{\prime}),\phi(s))$ in an attempt to estimate the performance contribution of a demonstration (Eq. 3). In our RoboMimic experiments (Fig. 3), this approach performs comparably to Demo-SCORE and, in some cases, even outperforms it—without requiring the training of any additional models. However, Success Similarity performs consistently worse than Cupid across all tasks, supporting prior findings that influence functions offer a substantially stronger causal signal than heuristic measures of similarity [2].
@@ D.1 Proof of Proposition 1
As presented in Section 3, applying the basic derivation of the influence function\@footnotemark in [13] gives us that
@@ D.2 Derivation of Performance Influence for Variable Length Trajectories
Assume that $\theta(\mathcal{D})=\arg\min_{\theta^{\prime}}\mathcal{L}_{\mathrm{bc}}^{\prime}(\theta^{\prime};\mathcal{D})$, that $\mathcal{L}_{\mathrm{bc}}^{\prime}$ is twice differentiable in $\theta$, and that $H_{\mathrm{bc}}\succ 0$ is positive definite (i.e., $\theta(\mathcal{D})$ is not a saddle point)\@footnotemark. Then, it holds that
@@
As presented in Section 3, applying the basic derivation of the influence function\@footnotemark in [13] gives us that

### v2

@@ Abstract
In robot imitation learning, policy performance is tightly coupled with the quality and composition of the demonstration data. Yet, developing a precise understanding of how individual demonstrations contribute to downstream outcomes—such as closed-loop task success or failure—remains a persistent challenge. We propose Cupid, a robot data curation method based on a novel influence function-theoretic formulation for imitation learning policies. Given a set of evaluation rollouts, Cupid estimates the influence of each training demonstration on the policy’s expected return. This enables ranking and selection of demonstrations according to their impact on the policy’s closed-loop performance. We use Cupid to curate data by 1) filtering out training demonstrations that harm policy performance and 2) subselecting newly collected trajectories that will most improve the policy. Extensive simulated and hardware experiments show that our approach consistently identifies which data drives test-time performance. For example, training with less than 33% of curated data can yield state-of-the-art diffusion policies on the simulated RoboMimic benchmark, with similar gains observed in hardware. Furthermore, hardware experiments show that our method can identify robust strategies under distribution shift, isolate spurious correlations, and even enhance the post-training of generalist robot policies. Videos and code are made available at: https://cupid-curation.github.io. 00footnotetext: Correspondence to: cagia@cs.stanford.edu
@@ 2 Related Work
Data Curation in Robotics. Assembling larger and more diverse datasets has been central to scaling efforts in robot imitation learning [7, 8, 17, 18, 19, 20, 21], yet how to extract greater utility from these datasets remains an open question. Several works have explored data augmentation [22, 23, 24, 25, 26] and mixture optimization [27]. Only recently has attention shifted to valuating individual demonstrations for data curation [9, 10, 11]. Hejna et al. [10] estimate demonstration quality offline via mutual information—without considering policy performance—while Chen et al. [11] train classifiers to distinguish successful and failed rollouts across policy checkpoints. In contrast, we directly measure the causal influence of each demonstration on the policy’s expected return, providing a signal that (a) does not require observing both successes and failures, (b) uses only a single policy checkpoint, (c) is robust to spurious correlations in the policy’s rollout distribution, and (d) naturally extends to selecting new data, whereas [10, 11] only filter existing data. Concurrent to our work is DataMIL [28], which uses datamodels to select from large multi-task datasets with an offline metric, whereas we focus on single-task curation with an influence measure that directly reflects closed-loop returns from online policy rollouts.
@@ 3 Background: Data Attribution via Influence Functions
where $H_{\theta}=\frac{1}{n}\sum_{i=1}^{n}\nabla_{\theta}^{2}\ell(z^{i};\theta(\mathcal{D}))$ denotes the Hessian of the training loss [^1] To reduce the computational cost of Eq. 1, we use TRAK [2], which leverages random projections and a Gauss-Newton Hessian approximation for efficient influence estimation. This also makes the influence function amenable to the non-smooth, non-convex loss functions in practical deep learning problems, so we assume Eq. 1 is well-defined throughout this paper. [13]. Therefore, we can use the influence function to directly approximate the leave-one-out influence $\Psi^{\mathrm{out}}$ of a sample $z^{i}\in\mathcal{D}$ as $\Psi^{\mathrm{out}}_{\mathrm{inf}}(\hat{z},z^{i}):=-\frac{1}{n}\Psi_{\mathrm{inf}}(\hat{z},z^{i})$. In addition, for $z\not\in\mathcal{D}$ we similarly define the add-one-in influence as $\Psi^{\mathrm{in}}_{\mathrm{inf}}(\hat{z},z):=\frac{1}{n}\Psi_{\mathrm{inf}}(\hat{z},z)\approx f(\hat{z};\theta(\mathcal{D}\cup\{z\}))-f(\hat{z};\theta(\mathcal{D}))$ with $z$ excluded from the Hessian $H_{\theta}$.
@@ 4 Problem Formulation
Let $\Xi^{+}_{k}=\left\{S\subseteq\mathcal{H}\;\middle|\;|S|=k\right\}$ denote all possible $k$-demonstration subsets of a holdout dataset $\mathcal{H}=\{\xi^{1},\ldots,\xi^{n^{\prime}}\}$, where $k\leq n^{\prime}$. Determine which $k$ demonstrations should be added to $\mathcal{D}$ from $\mathcal{H}$ to maximize policy performance with respect to the task objective $J$. That is, find
@@
Policy Testing & Evaluation: To make progress on 1 and 2, we assume access to a small dataset of $m$ rollouts $\mathcal{D}_{\tau}=\{\tau^{1},\ldots,\tau^{m}\}\overset{\textup{iid}}{\sim}p(\tau|\pi_{\theta})$ of the base policy $\pi_{\theta}$ along with their associated returns $\{R(\tau^{1}),\ldots,R(\tau^{m})\}$ to estimate $J(\pi_{\theta})$. This aligns with how we currently evaluate policies in practice [43], despite lacking principled strategies to leverage evaluations towards BC policy improvement.
@@ 5.1 Demonstration-Performance Influence
Assume that $\theta(\mathcal{D})=\arg\min_{\theta^{\prime}}\mathcal{L}_{\mathrm{bc}}(\theta^{\prime};\mathcal{D})$, that $\mathcal{L}_{\mathrm{bc}}$ is twice differentiable in $\theta$, and that $H_{\mathrm{bc}}\succ 0$ is positive definite (i.e., $\theta(\mathcal{D})$ is not a saddle point)1. Then, it holds that [^2] Note that the fraction $1/H$ appears from the assumption that all trajectories have equal length, which we make purely for notational simplicity without loss of generality. We refer to Section D.2 for the variable length case.
@@
Estimate $\Psi_{\pi\text{-}\mathrm{inf}}$: First, evaluate the policy $\pi_{\theta}$ online to gather a set of rollouts $\mathcal{D}_{\tau}=\{\tau^{1},\ldots,\tau^{m}\}\overset{\textup{iid}}{\sim}p(\tau|\pi_{\theta})$ and their associated returns $\{R(\tau^{1}),\ldots,R(\tau^{m})\}$. Then, construct an empirical estimate of the performance influence $\widehat{\Psi}_{\pi\text{-}\mathrm{inf}}$ using Eq. 3, by averaging action influences across the rollouts in $\mathcal{D}_{\tau}$.
@@
1Input: Policy $\pi_{\theta}$, training data $\mathcal{D}$, demonstration $\xi$, data attribution method $\Psi$
2 Collect rollouts $\mathcal{D}_{\tau}=\{\tau^{1},\ldots,\tau^{m}\}\overset{\textup{iid}}{\sim}p(\tau|\pi_{\theta})$ and returns $\{R(\tau^{1}),\ldots,R(\tau^{m})\}$
@@ 9 Limitations
Selection methods. While the greedy selection procedures used in Eq. 4 and Eq. 5 are tractable to optimize and often improve over quality- and similarity-based measures [3], they ignore the interactions between demonstrations in the curated set [14, 48]. This can temper performance gains when the size of the curated set is large. Future work should investigate higher-order approximations that consider the joint diversity of the curated dataset, as is common in the active learning literature (e.g., [52, Sec. 4.3]).
@@ Acknowledgments
The appendix offers additional details w.r.t. the implementation of Cupid (Appendix A), the experiments conducted (Appendix B), along with extended results and analysis (Appendix C), and finally, supporting derivations for our data curation methods (Appendix D). Videos and code are made available at: https://cupid-curation.github.io.
@@ A.1 Influence Functions for Diffusion Policies
Assume that $\theta(\mathcal{D})=\arg\min_{\theta^{\prime}}\mathcal{L}_{\mathrm{bc}}(\theta^{\prime};\mathcal{D})$, that $\mathcal{L}_{\mathrm{bc}}$ is twice differentiable in $\theta$, and that $H_{\mathrm{bc}}\succ 0$ is positive definite (i.e., $\theta(\mathcal{D})$ is not a saddle point)1. Then, it holds that
@@
[…]
@@
Influence Approximations: Since the denoising loss $\ell$ in Eq. 7 is proportionate to the variational lower bound on the action log-likelihood $\log\pi_{\theta}(a|s)$, it may seem intuitive to substitute $\nabla_{\theta}\log\pi_{\theta}(a^{\prime}|s^{\prime})$ with $\nabla_{\theta}\ell(s^{\prime},a^{\prime};\pi_{\theta})$—assuming gradient alignment—to approximate the action influence (Eq. 2) as
@@
[…]
@@
    Using $\ell_{\mathrm{square}}$ significantly reduces the computational cost of computing action influences for policies with high-dimensional action spaces, because the $\ell^{2}$-norm collapses the model’s prediction into a scalar $\epsilon_{\theta}(\sqrt{\bar{\alpha}_{i}}a+\sqrt{1-\bar{\alpha}_{i}}\epsilon^{i},s,i)||^{2}$. As a result, computing Eq. 9 requires only a single model gradient $\nabla_{\theta}\ell_{\mathrm{square}}$ per training and test sample. In contrast, while the technique proposed by Lin et al. [55] offers a more accurate estimate of the leave-one-out divergence $D_{\mathrm{KL}}(\pi_{\theta(\mathcal{D})}(a^{\prime}|s^{\prime})||\pi_{\theta(\mathcal{D}\setminus(s,a))}(a^{\prime}|s^{\prime}))$, its computational cost scales linearly with the dimensionality of the model’s output, which may be prohibitive.
@@ A.2 CUPID Hyperparameters
[…]
@@ B.4 Baseline Details
Comparison to Performance Influence (Cupid): One can interpret Success Similarity as replacing the action influence $\Psi_{a\text{-}\mathrm{inf}}((s^{\prime},a^{\prime}),(s,a))$ (Eq. 2) with a state-based proxy $D(\phi(s^{\prime}),\phi(s))$ in an attempt to estimate the performance contribution of a demonstration (Eq. 3). In our RoboMimic experiments (Fig. 3), this approach performs comparably to Demo-SCORE and, in some cases, even outperforms it—without requiring the training of any additional models. However, Success Similarity performs consistently worse than Cupid across all tasks, supporting prior findings that influence functions offer a substantially stronger causal signal than heuristic measures of similarity [2].
@@ D.1 Proof of Proposition 1
As presented in Section 3, applying the basic derivation of the influence function1 in [13] gives us that
@@ D.2 Derivation of Performance Influence for Variable Length Trajectories
Assume that $\theta(\mathcal{D})=\arg\min_{\theta^{\prime}}\mathcal{L}_{\mathrm{bc}}^{\prime}(\theta^{\prime};\mathcal{D})$, that $\mathcal{L}_{\mathrm{bc}}^{\prime}$ is twice differentiable in $\theta$, and that $H_{\mathrm{bc}}\succ 0$ is positive definite (i.e., $\theta(\mathcal{D})$ is not a saddle point)1. Then, it holds that
@@
As presented in Section 3, applying the basic derivation of the influence function1 in [13] gives us that